# HyperMARL: Adaptive Hypernetworks for Multi-Agent RL

### v2

@@ D.1.1 Training and Evaluation
    • For Dispersion (5.2), we run 5 seeds and train for 20 million timesteps.
@@
D.1.2 Measuring Policy Diversity Details
We measure team diversity using the System Neural Diversity (SND) metric (Equation 9 (Bettini et al., 2023), details Section A) with Jensen-Shannon distance. SND ranges from 0 (identical policies across all agents) to 1 (maximum diversity). We collect a dataset of observations from IPPO-NoPS and IPPO-FuPS policies checkpointed at 5 and 20 million training steps. Each policy is rolled out for 10,000 episodes, generating 16 million observations. We then sample 1 million observations from this dataset to calculate the SND for each method tested.
D.2 HyperMARL Architecture
For the experiments in Sections 5.2 and 5.2, we use feedforward architectures, while for Section 5.3, we use recurrent GRU networks (Cho, 2014). For recurrent PPO variants, HyperMARL maintains the same architecture as for feedforward methods, generating only the actor and critic feedforward weights, not the GRU weights.
Appendix E Detailed Results
E.1 Dispersion Detailed Results
E.1.1 Sample Efficiency Plots
(a) IPPO Sample Efficiency
(b) MAPPO Sample Efficiency
Figure 13: Sample efficiency of IPPO and MAPPO variants on the Dispersion environment. While FuPS methods struggle to converge to optimal performance, Hypernetworks achieve efficiency comparable to NoPS. Even when running MAPPO-FuPS for longer (see Figure 14), it converges to suboptimal performance.
@@ E.1.1 Sample Efficiency Plots
E.2 Detailed MAMujoco Plots
@@
Figure 15: Performance of Recurrent IPPO and MAPPO on SMAX. HyperMARL performs comparably to these baselines.
E.3 Detailed Navigation Plots
(a) Same Goal (8 agents)
(b) Different Goals (8 agents)
(c) Four Goals (8 agents)
(d) Scenarios with 8 Agents
(e) Same Goal (4 agents)
(f) Different Goals (4 agents)
(g) Two Goals (4 agents)
(h) Scenarios with 4 Agents
(i) Same Goal (2 agents)
(j) Different Goals (2 agents)
(k) Scenarios with 2 Agents
Figure 17: Sample Efficiency of IPPO variants in the VMAS Navigation environment under different goal configurations and agent numbers.
E.4 Additional Ablations
@@
Appendix F Hyperparameters
Table 8: Hyperparameters for Specialisation and Synchronisation Game
@@ Appendix F Hyperparameters
Number of agents and foods 2, 4, 8
Number of foods 3
Number of seeds 5
Number of trials 10,000
Evaluation trials 1,000
Batch size 10
Learning rate 0.01
@@
Total Number of Parameters
@@ 2 Agents
NoPS: 36
FuPS: 34
FuPS+ID: 42
@@ 4 Agents
NoPS: 112
FuPS: 100
FuPS+ID: 148
@@ 8 Agents
NoPS: 384
FuPS: 328
FuPS+ID: 552
@@
Table 17: Default algorithm and model hyperparameters for the Ant-v2-4x2 environment (from  (Zhong et al., 2024)).
@@
Table 18: Default algorithm and model hyperparameters for the Humanoid-v2-17x1 environment (from  (Zhong et al., 2024)).
@@
Table 19: Default algorithm and model hyperparameters for the Walker2d-v2-2x3 environment (from  (Zhong et al., 2024)).
@@
Table 20: Default algorithm and model hyperparameters for the HalfCheetah-v2-2x3 environment (from  (Zhong et al., 2024)).
@@
Parameter Humanoid-v2-17x1 Walker2d-v2-2x3 HalfCheetah-v2-2x3 Ant-v2-4x2 Sweeps
AGENT_ID_EMBEDDING_DIM 64 64 64 64 None
HYPERNET_HIDDEN_DIMS 64 64 64 64 None
@@
smacv2_10_units IPPO 0.005,0.001,0.0003,0.004 0.001 - -
MAPPO 0.002,0.005,0.0003 0.0003 - -
@@
smacv2_20_units IPPO 0.002,0.005,0.0003 0.005 - -
MAPPO 0.002,0.005,0.0003 0.0003 - -

### v3

@@ G.2.1 Training and Evaluation
    • For Dispersion (5.2), we run 10 seeds and train for 20 million timesteps.
@@
G.2.2 Measuring Policy Diversity Details
We measure team diversity using the System Neural Diversity (SND) metric (Equation 5 [7], details Section D) with Jensen-Shannon distance. SND ranges from 0 (identical policies across all agents) to 1 (maximum diversity). We collect a dataset of observations from IPPO-NoPS and IPPO-FuPS policies checkpointed at 5 and 20 million training steps. Each policy is rolled out for 10,000 episodes, generating 16 million observations. We then sample 1 million observations from this dataset to calculate the SND for each method tested.
Appendix H Detailed Results
H.1 Kaleidoscope Off-Policy Comparison Details
Our comparison with Kaleidoscope [26], mentioned in Section 5, is conducted using off-policy methods due to its original design. Kaleidoscope incorporates intricate mechanisms (e.g., learnable masks, an ensemble of five critics, a specific diversity loss) and numerous specialised hyperparameters (e.g., for critic ensembling: ‘critic_deque_len‘, ‘critic_div_coef‘, ‘reset_interval‘; for mask and threshold parameters: ‘n_masks‘, ‘threshold_init_scale‘, ‘threshold_init_bias‘, ‘weighted_masks‘, ‘sparsity_layer_weights‘, etc.). Porting this full complexity to an on-policy PPO backbone would constitute a significant research deviation rather than a direct benchmark of the established method.
Therefore, we utilised Kaleidoscope’s original off-policy implementation to ensure a meaningful comparison. We adopted MATD3 as the algorithmic backbone for this evaluation, as it was the only publicly available Kaleidoscope variant with tuned hyperparameters for Multi-Agent MuJoCo (MaMuJoCo). The MaMuJoCo tasks were chosen for alignment with our primary on-policy (IPPO) results and Kaleidoscope’s original evaluation, specifically: Ant-v2, HalfCheetah-v2, Walker2d-v2 (overlapping with our IPPO experiments), and Swimmer-v2-10x2 ( which represents the MaMuJoCo variant with the highest number of agents). Comparative results in Table 5 show that HyperMARL achieves competitive results with Kaleidoscope, while only using two critics (standard MATD3) and without additional diversity objectives.
Table 5: Mean episode return in MAMuJoCo for off-policy MATD3 variants. IQM of the mean episode returns with 95% stratified bootstrap CI. Bold indicates the highest IQM score; * indicates scores whose confidence intervals overlap with the highest. Although Kaleidoscope employs an ensemble of five critics and an explicit diversity loss, HyperMARL (using a standard MATD3 setup with two critics) achieved competitive results without these additional mechanisms.
Environment Ind. Actors (MATD3) HyperMARL (MATD3) Kaleidoscope (MATD3)
Ant-v2 5270.38 (4329.73, 5719.78) *5886.58 (5840.00, 5920.66) 6160.70 (5798.02, 6463.83)
HalfCheetah-v2 *6777.04 (3169.11, 8233.94) 7057.44 (3508.70, 8818.11) *6901.00 (3609.73, 8192.38)
Walker2d-v2 *5771.87 (5144.84, 8103.34) 7057.68 (5976.50, 8166.09) *6664.32 (5408.95, 8828.11)
Swimmer-v2-10x2 *453.74 (427.24, 487.86) 465.91 (410.82, 475.77) *462.48 (444.22, 475.64)
H.2 Dispersion Detailed Results
H.2.1 Interval Esimates Dispersion
(a) IPPO
(b) MAPPO
Figure 18: Performance of IPPO and MAPPO on Dispersion after 20 million timesteps. We show the Interquartile Mean (IQM) of the Mean Episode Return and the 95% Stratified Bootstrap Confidence Intervals (CI) using [1]. Hypernetworks achieve comparable performance to NoPS, while FuPS struggle with specialisation.
@@ H.2.1 Interval Esimates Dispersion
H.3 Detailed MAMujoco Plots
@@
Figure 20: Performance of Recurrent IPPO and MAPPO on MaMoJoCo. HyperMARL performs comparably to these baselines, and is the only method with shared actors to demonstrate stable learning in the notoriously difficult 17-agent Humanoid environment.
H.4 Detailed Navigation Plots
(a) 8 Agents, Same Goal
(b) 8 Agents, Different Goals
(c) 8 Agents, Four Goals
(d) 4 Agents, Same Goal
(e) 4 Agents, Different Goals
(f) 4 Agents, Two Goals
(g) 2 Agents, Same Goal
(h) 2 Agents, Different Goals
Figure 21: Sample efficiency of IPPO variants in the VMAS Navigation environment. Plots show IQM and 95% CI (shaded regions) of mean episode return against training steps for different agent counts (rows: 8, 4, 2 agents) and goal configurations (columns, where applicable: Same, Different, Specific Goal Counts). Legend shown at bottom applies to all subplots.
H.5 Interval Esimates - SMAX
(a) 2s3z
(b) 3s5z
(c) SMACv2 10 Units
(d) SMACv2 20 Units
Figure 22: Performance comparison in SMAX environments after 10 million timesteps. We show the Interquartile Mean (IQM) of the Mean Win Rate and the 95% Stratified Bootstrap Confidence Intervals (CI). HyperMARL performs comparably to FuPS baselines across all environments, demonstrating its effectiveness in tasks requiring homogeneous behaviours and using recurrency.
H.6 Additional Ablations
@@
Appendix I Hyperparameters
Table 6: Hyperparameters, Training and Evaluation for Specialisation and Synchronisation Game
@@ Appendix I Hyperparameters
Environment Configuration
Number of agents 2, 4, 8, 16
Maximum steps per episode 10
Training Protocol
Number of seeds 10
Training steps 10,000
Evaluation episodes 100
Evaluation interval 1,000 steps
Batch size 32
Model Architecture
Hidden layer size 8, 16, 32, 64
@@
Model Parameter Count
@@ 2 Agents
NoPS: 60
FuPS: 58
FuPS+ID: 74
FuPS+ID (No State): 42
@@ 4 Agents
NoPS: 352
FuPS: 240
FuPS+ID: 404
FuPS+ID (No State): 148
@@ 8 Agents
NoPS: 2400
FuPS: 2344
FuPS+ID: 2600
FuPS+ID (No State): 552
16 Agents
NoPS: 17728
FuPS: 17488
FuPS+ID: 18512
FuPS+ID (No State): 2128
@@ 16 Agents
Table 15: Default algorithm and model hyperparameters for the Ant-v2-4x2 environment (from  [49]).
@@
Table 16: Default algorithm and model hyperparameters for the Humanoid-v2-17x1 environment (from  [49]).
@@
Table 17: Default algorithm and model hyperparameters for the Walker2d-v2-2x3 environment (from  [49]).
@@
Table 18: Default algorithm and model hyperparameters for the HalfCheetah-v2-2x3 environment (from  [49]).
@@
Parameter Humanoid Walker2d HalfCheetah Ant Sweeps
v2-17x1 v2-2x3 v2-2x3 v2-4x2
AGENT_ID_DIM 64 64 64 64 None
HNET_HIDDEN_DIMS 64 64 64 64 None
@@
smacv2_10_units IPPO 0.005, 0.001, 0.0003, 0.004 0.001
MAPPO 0.002, 0.005, 0.0003 0.0003
@@
smacv2_20_units IPPO 0.002, 0.005, 0.0003 0.005
MAPPO 0.002, 0.005, 0.0003 0.0003
@@
Appendix J Computational Resources
Table 22: Computational Resources by Experiment Type
Experiment Category Hardware Configuration Execution Time Total Hours
Specialisation, 8 cores on AMD EPYC 7H12 2-6 hours per run 250
Synchronisation & Dispersion 64-Core Processor (agent-count dependent)
Navigation 8 cores on AMD EPYC 7H12 4-10 hours per run 320
Experiments 64-Core Processor + NVIDIA RTX A4500
MaMuJoCo 8 cores on AMD EPYC 7H12 8-24 hours per run 1,680
Experiments 64-Core Processor + NVIDIA RTX A4500 (scenario & algorithm dependent)
SMAX 8 cores on AMD EPYC 7H12 2-5 hours per run 160
Experiments 64-Core Processor + NVIDIA RTX A4500